### v1

# ExoIris: fast exoplanet transmission spectroscopy in Python

Abstract
I present ExoIris, a user-friendly Python package for exoplanet transmission and emission spectroscopy. Unlike existing tools, ExoIris models two-dimensional spectrophotometric transit time series directly and supports the joint analysis of multiple datasets obtained with different instruments and at different epochs. These features enable a self-consistent estimation of both wavelength-independent and wavelength-dependent parameters. They offer a more robust workflow compared to the commonly used two-step approach, where a ”white” light curve is fitted first and the transmission spectrum is then derived from independent fits constrained by the white-light solution. Despite its increased flexibility and robustness, ExoIris remains computationally efficient. A low-resolution transmission spectrum can be estimated from a single JWST NIRISS transit observation in $\sim 5$ minutes assuming white noise, and in $\sim 15$ minutes when accounting for time-correlated systematics using a Gaussian process noise model, on a standard desktop computer.
Exoplanets (498) — Transmission spectroscopy (2133) — Open source software (1866)
software: PyTransit (Parviainen, 2015; Parviainen & Korth, 2020; Parviainen, 2020), LDTk (Parviainen & Aigrain, 2015), Astropy (Robitaille et al., 2013; Collaboration et al., 2018, 2022) emcee (Foreman-Mackey et al., 2013), celerite2 (Foreman-Mackey et al., 2023), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Numba (Lam et al., 2015).
show]hannu@iac.es
1 Introduction
I present ExoIris, [^1] The ExoIris package is available via PyPI and GitHub (https://github.com/hpparvi/ExoIris). Documentation and tutorials can be found at https://exoiris.readthedocs.io. an open-source Python package for transmission and emission spectroscopy of exoplanets. ExoIris differs from existing tools by directly modeling two-dimensional spectrophotometric transit datasets (time series of spectroscopic transit or eclipse observations) and by supporting joint analyses of multiple datasets obtained with different instruments and at different epochs.
In ExoIris, wavelength-dependent quantities (planet-star radius ratio and stellar limb darkening profile) and wavelength-independent quantities (e.g., orbital elements) are inferred jointly from the full dataset. This contrasts with the common two-step approach, where each spectroscopic light curve is fitted independently, often using tight priors from a white-light curve analysis and fixed limb darkening coefficients from stellar atmosphere models. Such methods fail to use the information content of the observations fully and may underestimate uncertainties. By modeling the full two-dimensional dataset, ExoIris enables self-consistent inference of all relevant parameters, which results in more robust parameter estimates and transmission spectra.
ExoIris also supports joint modeling of heterogeneous datasets, allowing users to account for instrument- and time-dependent effects such as transit timing variations (TTVs), instrumental systematics (e.g. Moran et al., 2023; Alderson et al., 2025), and stellar contamination (e.g. Rackham et al., 2018, 2019; Rackham & de Wit, 2024). Time-correlated noise and systematics can be modeled with Gaussian processes (GPs), using fixed or inferred hyperparameters.
Despite its flexibility, ExoIris remains computationally efficient. The planet-star radius ratio is modeled as a spline over wavelength, with freely located knots treated as free parameters. This decouples the spectral resolution of the transmission spectrum from that of the data, enabling high-resolution sampling of sharp absorption features while allowing the smoother continuum to be modeled at lower resolution. The number of free parameters in the final model is primarily set by the number of radius ratio spline knots, and adding new datasets does not significantly increase model dimensionality if the wavelength space is already covered. For example, a joint fit to a single JWST NIRISS transit with $\mathcal{R}\sim 300$, 100 radius ratio knots, and a GP noise model typically completes in 10–20 minutes on a standard desktop computer. More complex analyses, such as the joint modeling of eight WASP-39 b Early Release Science datasets (Sect. 3), complete within a few hours.
2 ExoIris
2.1 Spectrophotometry Model Overview
An ExoIris transmission spectroscopy analysis considers a collection of spectrophotometric datasets, where each dataset contains spectrophotometric fluxes and uncertainties from a single transit observation, along with time and wavelength sampling and parameter grouping metadata. The observed spectrophotometric flux for dataset $d$, as a function of time $t$ and wavelength $\lambda$, is modeled as
$F_{\mathrm{d}}(t,\lambda)=O+(1-O)B_{\mathrm{d}}(\lambda)\;T\left(t,k(\lambda),\mathbb{l}(\lambda),\boldsymbol{\Theta}\right),$ (1)
where $O$ is a flux offset term, $B$ is a baseline flux model, and $T$ is the spectrophotometric transit model. The model depends on the planet-star radius ratio $k$, the limb darkening parameters $\mathbb{l}$, and the orbital parameter vector $\boldsymbol{\Theta}$ (comprising the zero epoch $t_{0}$ for each epoch group, orbital period $p$, inclination $i$, eccentricity $e$, and argument of periastron $\omega$). Each dataset is assigned to offset, noise, epoch, and baseline model groups, which determine the model parameterization. These are described in detail in Sects. 2.5, 2.8, 2.7, and 2.6, respectively.
Figure 1: Example of a spectrophotometric transit model. The first two panels show the power-2 limb darkening coefficients as splines over wavelength with freely placed knots. The third panel shows the planet’s radius ratio modeled in the same way. The rightmost panel displays the model flux as a function of time and wavelength.
The planet-star radius ratio, $k$, is modeled as a spline over wavelength, decoupling the resolution of the transmission spectrum from the spectroscopic resolution of the data (Sect. 2.3). The limb darkening vector $\mathbb{l}$ depends on the chosen model. Currently, ExoIris supports two-parameter analytical limb darkening laws with coefficients modeled as splines and a numerical LDTkLDM-based model (Sect. 2.4). In the analytical case, $\mathbb{l}$ consists of limb darkening coefficients at the knot wavelengths; in the numerical case, it contains the stellar effective temperature, metallicity, and surface gravity.
Figure 1 illustrates a single spectrophotometric model for a dataset with 1000 exposures and 600 wavelength elements. The model flexibly captures both sharp and smooth variations in $k$, allowing high resolution where needed and lower resolution elsewhere.
2.2 Transmission Spectroscopy Transit Model
ExoIris uses the TSModel transit model from PyTransit v2.6 (Parviainen, 2015; Parviainen & Korth, 2020; Parviainen, 2020), specifically designed for transmission spectroscopy. It is based on the latest version of the RoadRunner model (Parviainen, 2020), which enables fast light curve evaluation with arbitrary radially symmetric stellar surface brightness profiles. This flexibility supports both analytical limb darkening laws and numerical profiles derived from stellar atmosphere models.
Assuming a radially symmetric stellar intensity profile $I(z)$, where $z$ is the projected distance from the stellar center normalized to the stellar radius, the in-transit flux is given by
$I$ $=\int_{A_{\star}}I(z)\;\mathrm{d}A-\int_{A_{\mathrm{p}}}I(z)\;\mathrm{d}A$ (2)
$=2\pi\int_{0}^{1}z\;I(z)\;\mathrm{d}z-\int_{A_{\mathrm{p}}}I(z)\;\mathrm{d}A$ (3)
$=\hat{I}_{\star}A_{\star}-\hat{I}_{\mathrm{p}}A_{\mathrm{p}},$ (4)
where $\hat{I}_{\star}$ is the mean surface brightness of the stellar disk, $A_{\star}$ its area, $\hat{I}_{\mathrm{p}}$ the mean surface brightness of the stellar disk occulted by the planet, and $A_{\mathrm{p}}$ the occulted area. If the stellar radius and out-of-transit flux are normalized to unity, the expression simplifies to
$I=\frac{\pi\hat{I}_{\star}-\hat{I}_{\mathrm{p}}A_{\mathrm{p}}}{\pi\hat{I}_{\star}}=\frac{I_{\star}-\hat{I}_{\mathrm{p}}A_{\mathrm{p}}}{I_{\star}},$ (5)
where $I_{\star}=\pi\hat{I}_{\star}$ is the total stellar flux.
Among these quantities, $I_{\star}$ and $A_{\mathrm{p}}$ are straightforward to calculate. The main computational challenge lies in evaluating $\hat{I}_{\mathrm{p}}$ efficiently for arbitrary limb darkening profiles. The RoadRunner model addresses this by expressing $\hat{I}_{\mathrm{p}}$ as a weighted sum over the stellar intensity profile, reducing the model evaluation to a dot product between two one-dimensional arrays multiplied by the area occulted by the planet:
$I=\frac{I_{\star}-\mathbf{w}(k,b)\cdot\mathbf{l}\times a(k,b)}{I_{\star}},$ (6)
where $\mathbf{w}(k,b)$ is a weight vector dependent on the radius ratio $k$ and sky-projected planet-star center separation $b$ as described in Parviainen (2020), $\mathbf{l}$ is the discretised limb darkening profile, and $a(k,b)$ is the occulted area.
The mean occulted brightness $\mathbf{w}\cdot\mathbf{l}$ varies smoothly with $b$ for fixed $k$, and can be precomputed as a one-dimensional interpolation table to further speed up evaluation. The resulting model becomes
$I=\frac{I_{\star}-\hat{I}_{\mathrm{p}}(b)\times a(k,b)}{I_{\star}},$ (7)
where $\hat{I}_{\mathrm{p}}(b)$ is the precomputed mean occulted intensity for fixed $k$, interpolated as a function of $b$. The intersection area $a(k,b)$ is computed using the circle-circle intersection routine by Agol et al. (2020), which also provides its derivatives with respect to $k$ and $b$.
Separating the limb darkening term from the geometric occultation enables high computational efficiency, making the RoadRunner model well suited for transmission spectroscopy. A transmission spectroscopy dataset derived from a spectroscopic time series consists of flux estimates in multiple narrow passbands, all sampled at identical points in time. This allows for several optimizations that significantly reduce the computational cost of the model:
  1. 1.
    The planet-star center distance, $b$, is independent of passband and only needs to be computed once per exposure.
  2. 2.
    Although the weight vector $\mathbf{w}$ depends on $k$, the derivative $\partial\hat{I}_{\mathrm{p}}/\partial k$ is negligible compared to other effects of $k$ variation. Thus, the weight vector computed for the mean radius ratio $\hat{k}$ can be used across all passbands with minimal impact on accuracy.
  3. 3.
    While computing the circle-circle intersection area is already inexpensive, it can be further accelerated. In the fully occulted regime ($b<1-\max(\mathbf{k})$, where $\mathbf{k}$ is the vector of radius ratios across passbands), the intersection area simplifies to
    $a(k,b)=\pi k^{2}.$ (8)
    In the partial overlap regime ($1-\max(\mathbf{k})\leq b\leq 1+\max(\mathbf{k})$), a first-order Taylor expansion around $\hat{k}$ gives
    $a(k,b)=a(\hat{k},b)+(k-\hat{k})\frac{\partial a}{\partial k}$ (9)
[…]
The last approximation introduces minor errors during ingress and egress (1k<b<1+k1-k<b<1+k), but its effect on inferred transit depths and orbital parameters is negligible for realistic $k$ variations. Numerical tests [^2] https://github.com/hpparvi/PyTransit/blob/master/notebooks/tsmodel/test_accuracy.ipynb show that for a hot Jupiter like WASP-39 b ($k$ varying from 0.144 to 0.154), the maximum transit depth error is 0.01% (3 ppm), and for a sub-Neptune like TOI-836 c ($k$ from 0.033 to 0.036), the error is below 0.004% (0.05 ppm). For grazing hot Jupiters, the error can rise to $\sim$0.5%, but this remains below current observational precision and can be corrected if needed.
2.3 Radius Ratio Parameterization
In transmission spectroscopy, the key quantity of interest is the wavelength-dependent planet-star radius ratio, $k(\lambda)$, which encodes the shape of the transmission spectrum. While the TSModel enables the efficient calculation of thousands of spectrophotometric transit light curves, treating the radius ratio in each passband as an independent free model parameter would result in exceedingly large-dimensional models.
ExoIris addresses this by decoupling the data resolution from the spectral resolution of the retrieved transmission spectrum. As shown in Fig. 1, the radius ratio is modeled as a spline over wavelength, defined by a user-specified set of knots. The values at the knot positions are treated as free parameters, and the radius ratio for each passband is interpolated from the spline. Both the number and placement of the knots are fully customizable. In typical use, knots can be distributed uniformly or logarithmically in wavelength, or clustered around known or suspected spectral features. Although the transit model is evaluated at the full spectroscopic data resolution, the radius ratio values are constrained by the lower-dimensional spline.
This parameterization offers several benefits. First, model complexity depends on the number of spline knots rather than the data resolution. Second, the spectral resolution of the transmission spectrum can be adapted to the science case—for example, by placing more knots near expected absorption features and fewer in smooth continuum regions. Finally, the knot configuration can be modified dynamically during analysis, allowing for iterative refinement: an initial coarse (low-resolution) spectrum can be followed by a higher-resolution fit using the earlier result as a starting point.
2.4 Limb Darkening
2.4.1 Analytical Limb Darkening Models
ExoIris models the wavelength-dependent stellar limb darkening coefficients using the same spline-based approach as for the planet-star radius ratio: a user-defined set of wavelength knots defines a spline, and the coefficient values at these knots are treated as free parameters (see Fig. 1). The limb darkening coefficients for individual passbands are then obtained by interpolation.
This parameterization is available for all two-parameter analytical limb darkening laws supported by the RoadRunner model, including the quadratic and power-2 laws. It provides the same flexibility and regularization benefits as the spline-based radius ratio model: the number and placement of knots can be adjusted to balance spectral resolution against model complexity, and the knot configuration can be updated at any point during the analysis.
As with the radius ratio, users can increase the knot density in regions where limb darkening is expected to change rapidly, or reduce the number of free parameters in smoother spectral regions. When applied to a joint fit of all spectrophotometric light curves, this approach reduces reliance on external priors and allows limb darkening constraints to be derived empirically from the data. Although ExoIris supports the use of priors generated with the LDTk package (Parviainen & Aigrain, 2015)–which uses PHOENIX stellar atmosphere models (Husser et al., 2013) to generate theoretical limb darkening profiles–these priors can be made broad enough to ensure that the inference is driven by the data, not the theory, when discrepancies arise.
2.4.2 LDTk-Based Limb Darkening Model
In addition to the spline-based analytical models, ExoIris supports a numerical limb darkening model based on LDTk. This method uses PHOENIX stellar atmosphere models (Husser et al., 2013) to generate intensity profiles directly from three stellar parameters: effective temperature ($T_{\mathrm{eff}}$), surface gravity ($\log g$), and metallicity ($\mathrm{Fe}/\mathrm{H}]$). These parameters are treated as global free parameters, independent of the number of passbands, providing a low-dimensional and physically motivated alternative to spline interpolation.
While the LDTk-based model is more tightly constrained by stellar theory, the spline-based approach offers greater flexibility when empirical, data-driven constraints are preferred. Both methods are fully supported in ExoIris and can be chosen based on the desired balance between theoretical and observational constraints.
2.5 Additive Flux Offset
The offset term, $O$, in Eq. (1) accounts for additive bias-level variations arising from instrumental or astrophysical sources. These variations can be instrument- or epoch-dependent and directly affect the measured transit depths.
First, bias estimation errors in the instrument can lead to apparent depth differences between detectors, as seen in the NIRSpec observations of TOI-776 b (Moran et al., 2023; Alderson et al., 2025), where nonphysical jumps were observed between the transmission spectra from the NRS1 and NRS2 detectors. Second, the presence of unocculted stellar inhomogeneities, such as faculae or star spots, can cause wavelength-independent, epoch-dependent variations in the observed flux. This so-called ”transit light source effect” (Rackham et al., 2018, 2019; Rackham & de Wit, 2024) can imprint spectroscopic features that mimic atmospheric signals. While the effect is inherently wavelength-dependent, its relative variation across epochs can be approximated (to first order) as a constant offset in flux. Third, variable third-light contamination from an unresolved background source (e.g., a nearby variable star) can similarly produce epoch-dependent flux offsets.
ExoIris addresses these effects using an offset parameter, $O$, which modifies the light curve model in Eq. (1). Each dataset is assigned to an offset group, and a separate $O_{i}$ parameter is introduced for each group. Offset group 0 is the reference and has its offset fixed to zero, while the offsets for other groups are treated as free parameters. The offset $O$ affects the overall transit depth uniformly across all wavelengths within a group, leaving the relative shape of the transmission spectrum unchanged.
This implementation allows the model to absorb detector- or epoch-specific flux shifts without distorting the spectroscopic features of interest. It also provides a useful diagnostic: non-zero offsets may indicate instrumental or astrophysical systematics at the group level. However, the method relies on sufficient wavelength overlap between offset groups—either directly with the reference group or indirectly through overlapping intermediate groups. Without such overlap, the offset becomes degenerate with the radius ratio, and additional constraints are needed to resolve the degeneracy.
Figure 2 illustrates this functionality using a simplified version of the TOI-776 b dataset from Alderson et al. (2025), where two transits were observed with both the NRS1 and NRS2 detectors. In the second transit, a shift is visible between the NRS1 and NRS2 light curves, while no such shift appears in the first transit. Datasets 0 and 1 correspond to the first epoch and are assigned to the reference offset group (group 0). Datasets 2 and 3, from the second epoch, are assigned to offset groups 1 and 2, respectively, introducing two additional offset parameters. In the example shown, the model is evaluated with offset values of 0.0 and 0.5 for groups 1 and 2, resulting in identical flux models for datasets 0 and 2. However, the apparent transit depths for dataset 3 are reduced by 50% relative to dataset 1, even though the physical radius ratio remains unchanged.
Figure 2: Example illustrating the use of offset groups to account for instrument- or epoch-dependent bias-level variations arising from bias estimation errors or unocculted stellar inhomogeneities. The figure shows four spectrophotometric datasets from two transit epochs, each observed with two detectors. The first two columns show the spectrophotometric model for all datasets, the next two columns display the average light curves, and the final column shows the radius ratio as a function of wavelength. Here, DS is the dataset index, BG the offset group index, and EG the epoch group index (see Sect. 2.7).
2.6 Multiplicative Flux Baseline
In addition to the additive offset term, the flux model includes a wavelength-dependent multiplicative baseline term, $B_{\mathrm{d}}(\lambda)$, which represents the out-of-transit flux level. As with the radius ratio and limb darkening coefficients, the baseline is modeled as a spline over wavelength. However, each dataset is assigned its own independent baseline model, with a customizable number of spline knots (see Fig. 3).
The baseline model is not intended to capture short-scale spectral features—though this can be approximated by increasing the number of knots—nor is it designed to account for time-dependent variability, which is more appropriately handled by the GP noise model (Sect. 2.8). Instead, its primary purpose is to propagate uncertainties in the baseline flux level into the inferred transmission spectrum, ensuring that these uncertainties are properly reflected in the final parameter estimates.
Figure 3: Example of a baseline flux model in an ExoIris analysis with two spectrophotometric datasets. Each dataset is associated with a multiplicative baseline model, represented as a spline over wavelength. The spline knot values are included as free parameters in the model.
2.7 Transit Timing Variations
Transit timing variations (deviations from strict periodicity in the times of transit centers, TTVs) are typically caused by gravitational interactions between planets in multi-planet systems. While TTVs are not an issue when analyzing transits individually, they must be accounted for in joint analyses of observations from multiple epochs. Failing to do so can distort the light curve shape and bias the inferred planet-star radius ratio, leading to inaccurate transmission spectra.
ExoIris accounts for TTVs by assigning each dataset to an epoch group. Each epoch group introduces a separate transit center time as a free model parameter. This allows datasets to be grouped by observation epoch, with TTVs automatically incorporated into the model.
Figure 4 shows an example with four spectrophotometric datasets observed over three epochs. The first two datasets correspond to the same transit observed with different instruments, while the remaining two are from separate epochs and observed with a third instrument. In this setup, three transit center parameters are included in the model, one for each unique epoch index.
If TTVs are expected to be negligible, all datasets can be assigned to epoch group 0. However, assigning each transit epoch its own group adds minimal model complexity and can provide useful constraints. For high-precision datasets, the resulting transit center posteriors may reveal subtle TTV signals or help refine orbital parameters.
Figure 4: Example illustrating the use of epoch groups (EG) to account for transit timing variations in a joint transmission spectrum analysis with ExoIris. The example includes four spectrophotometric datasets: two from the same epoch observed with different instruments, and two from separate epochs observed with a third instrument.
2.8 Noise Modeling
Robust noise modeling is important in transmission spectroscopy, particularly when combining observations from multiple epochs or instruments. ExoIris provides a flexible approach to noise modeling based on the concept of noise groups, similar to the offset and epoch groups described earlier. Each spectrophotometric dataset is assigned to a specific noise group, allowing for tailored noise treatment across heterogeneous datasets.
ExoIris currently supports three noise modeling options: white noise, GPs with fixed hyperparameters, and GPs with free hyperparameters. GPs are implemented using the celerite2 package (Foreman-Mackey et al., 2023).
The white noise model assumes independent, normally distributed uncertainties, scaled by a single multiplicative factor per noise group. These uncertainties are provided as an array with the same dimensions as the flux array, containing one value per flux measurement. They can be supplied by the data reduction pipeline or estimated during dataset initialization. The noise scaling factor is included as a free parameter to account for possible under- or overestimation of the input uncertainties. This model is appropriate when the noise is dominated by photon noise and there is minimal time-correlated or systematic variability.
The GP model with fixed hyperparameters uses a user-defined kernel to describe correlated noise or long-term trends. Any kernel supported by celerite2 can be used, and ExoIris includes methods for optimizing hyperparameters for the Matern-3/2 kernel. This approach is suitable in most cases where correlated noise or longer-time-scale trends are present, and it offers a good balance between flexibility and computational efficiency.
The GP model with free hyperparameters allows the hyperparameters to be estimated jointly with the rest of the model parameters. Currently, this is supported only for the basic Matern-3/2 kernel and comes with a higher computational cost. Support for additional kernels may be added in future versions.
Although a single ExoIris analysis cannot combine different noise models, individual noise groups can use different GP kernels or kernel parameters. Additionally, users can change the noise model during an analysis or continue from a saved solution. This allows, for example, a quick initial fit using a white noise model to be followed by a more detailed analysis with a GP noise model.
2.9 Masking of Bad Data
Each ExoIris dataset includes a mask array that flags spectrophotometric data points to be excluded from the log-likelihood evaluation. Non-finite values, such as NaNs and infinities, in the flux or uncertainty arrays are automatically masked. In addition, ExoIris provides tools to detect and mask outliers, allowing users to remove problematic data points before or during the analysis.
3 Example: WASP-39 b Early Release Science Joint Analysis
The ExoIris example GitHub repository [^3] https://github.com/hpparvi/exoiris_examples contains a set of examples starting from simple analyses with one dataset and continuing to more complex cases combining many datasets into a joint analysis. The WASP-39 b Early Release Science (ERS) joint analysis provides a comprehensive demonstration of a joint transmission spectrum analysis using spectrophotometric light curves from WASP-39 b transits observed with NIRSpec, NIRSpec PRISM, NIRCam, and NIRISS. The data used in this example were produced by Carter et al. (2024), who reprocessed the original observations presented by Ahrer et al. (2023), Feinstein et al. (2023), Alderson et al. (2023), and Rustamkulov et al. (2023). The example illustrates how to configure bias, ephemeris, and noise groups for joint modeling and includes individual transmission spectroscopy analyses for each instrument.
3.1 Analysis Setup
Figure 5: ExoIris setup for the WASP-39 b ERS joint analysis. The top row shows the radius ratio knot locations, the second row shows the limb darkening knot locations, and the remaining rows show the central wavelengths of the spectroscopic elements for each dataset. The noise (N), offset (O), and epoch (E) group assignments are indicated beside each dataset’s wavelength coverage. Details on cropping and binning are provided in the text.
Figure 6: Spectrophotometric light curves used in the WASP-39 b joint analysis. White pixels indicate outliers that were masked from the analysis.
The joint analysis includes eight spectrophotometric datasets, as shown in Fig. 5. The NIRSpec PRISM data are cropped to $2.0\leq\lambda\leq 5.5$ $\mu\text{m}$ to remove saturated wavelengths (Carter et al., 2024; Rustamkulov et al., 2023), and then binned in time to 30 s (but not binned in wavelength). The NIRCam F322W2 data are cropped to $\lambda<4$ $\mu\text{m}$ and $t>2459783.32$, and binned in wavelength to a resolution of $R\approx 600$. The NIRSpec G395H data are split into two time segments ($2459791.45>t>2459791.62$ and $t>2459791.625$), cropped to $\lambda>2.85$ $\mu\text{m}$, and binned in wavelength to $R\approx 600$, resulting in four datasets. The NIRISS SOSS order 1 data are cropped to $0.86\leq\lambda\leq 2.8$ $\mu\text{m}$ and binned to $R\approx 300$, while order 2 data are cropped to $0.60\leq\lambda\leq 0.84$ $\mu\text{m}$ and binned to $R\approx 800$. Instead of using uncertainties from the reduction pipelines, we estimate a white noise scatter (standard deviation) for each light curve directly from the data.
Each of the four transit observations is assigned to a separate epoch group, adding four transit center parameters to the model. Although WASP-39 b shows no signs of TTVs, the Carter et al. (2024) NIRISS reduction produces a significant offset in transit center time relative to a linear ephemeris, in contrast to the other datasets (see Fig. 10). While this offset is a data reduction artifact, it serves to demonstrate the functionality of epoch groups in ExoIris. Each detector is assigned a separate bias offset group, resulting in five free offset parameters. The NIRSpec PRISM dataset is used as the reference (offset group 0) due to its wide wavelength coverage. Each dataset is also assigned its own noise group, resulting in six white noise scaling parameters.
The transmission spectrum is modeled using 126 radius ratio knots across three wavelength regions: 0.6–0.75 $\mu\text{m}$ with a spacing of approximately 0.02 $\mu\text{m}$, a denser region around the 0.768 $\mu\text{m}$ K line with a spacing of 0.004 $\mu\text{m}$, and 0.78–5.5 $\mu\text{m}$ with a spacing of approximately 0.04 $\mu\text{m}$. Limb darkening is modeled with 10 knots linearly spaced from 0.6 to 5.5 $\mu\text{m}$.
3.2 Results
Figure 7 shows a joint fit to the ”white” light curves. This is an optional initial step that can be used to initialize the transmission spectroscopy analysis, and is also useful for visual identification of spot crossings. Figure 8 shows the transmission spectrum estimated from the joint analysis, together with the transmission spectrum from Carter et al. (2024). Finally, Fig. 9 shows the residuals relative to the median posterior model, and Fig. 10 shows the posterior distributions for a set of quantities that are useful when assessing the quality of the observations and the analysis.
The ExoIris transmission spectrum agrees closely with the Carter et al. (2024) results with fitted limb darkening coefficients.
As mentioned in Carter et al. (2024), JWST observations offer an unprecedented precision for transit-based stellar density estimation (assuming the eccentricity and argument of periastron are also known to a high precision.) Joint modeling of multiple spectrophotometric data sets naturally increases precision beyond that of individual analyses. Assuming a circular orbit, the stellar density precision for WASP-39 obtained from the joint model is 0.2%.
Figure 7: WASP-39 b white light curves jointly fitted with ExoIris. Blue and yellow points show the observed photometry, and the black line shows the median posterior model. The NIRSpec NRS1 and NRS2 light curves are split in time due to a jump near mid-transit.
Figure 8: WASP-39 b transmission spectrum estimated from the joint analysis of NIRSpec, NIRSpec PRISM, NIRCam, and NIRISS data using ExoIris. The black line shows the median of the posterior model, blue shading indicates the 16th and 84th percentiles, gray points show the radius ratio values from Carter et al. (2024), and the orange line shows the Carter et al. (2024) values binned to a similar resolution with the analysis.
Figure 9: Residuals relative to the median posterior model for the WASP-39 b joint analysis.
Figure 10: Posterior distributions from the WASP-39 b joint analysis for transit centers, bias offsets, white noise scaling factors, and out-of-transit baseline levels. Transit centers are shown relative to the best-fit linear ephemeris. The NIRISS dataset shows a significant discrepancy due to reduction artifacts in Carter et al. (2024), not physical TTVs. All datasets show small bias offsets relative to the PRISM data.
4 Discussion
The ExoIris package provides a flexible and efficient framework for analyzing spectrophotometric time-series data in the context of exoplanet transmission spectroscopy. Its core design principles (direct modeling of two-dimensional spectrophotometric datasets and joint analysis of multi-instrument, multi-epoch observations) enable robust and self-consistent inference of both wavelength-dependent and wavelength-independent parameters.
A key strength of ExoIris lies in its explicit parameterization of wavelength-dependent quantities, such as the planet-star radius ratio and limb darkening coefficients, using spline models. This approach decouples the spectral resolution of the transmission spectrum from the resolution of the data, allowing users to balance model complexity against scientific goals. In addition, the inclusion of group-based structures for instrumental offsets, noise properties, and transit timing variations enables accurate joint modeling of heterogeneous datasets.
The example analysis of WASP-39 b highlights the effectiveness of these methods. ExoIris accommodates diverse instrument modes, detector configurations, and data characteristics within a single modeling framework. The results reproduce the transmission spectrum derived by Carter et al. (2024) while also providing improved flexibility for sensitivity testing, including the impact of systematic offsets and timing discrepancies.
I acknowledge support by the Spanish Ministry of Science and Innovation with the Ramon y Cajal fellowship number RYC2021-031798-I. Funding from the University of La Laguna and the Spanish Ministry of Universities is acknowledged.